OpenStudy (anonymous):

integrate t-3 from 0 to x^2

OpenStudy (anonymous):

\[\int\limits_{0}^{x^2} (t-3) dt\]

OpenStudy (anonymous):

Expand the integrand and write it like this \[\Large \int_0^{x^2}tdt-3 \int_0^{x^2}1dt \]

OpenStudy (anonymous):

then just apply regular exponential rules of integration. \[ \Large \int x^n dx= \frac{x^{n+1}}{n+1} \]

OpenStudy (anonymous):

okay, thanks . so what would be the answer, i'm trying to check my work.

OpenStudy (anonymous):

\[\Large \frac{t^2}{2}-3t \] Now just substitute the upper bounds \(t=x^2\)

OpenStudy (anonymous):

\[\Large \frac{x^4}{2}-3x^2 \]

OpenStudy (anonymous):

so our teacher was wrong ... thanks, i was wondering why i got a different answer.

OpenStudy (anonymous):

i already used that, i thought i might have entered something in wrong. thanks again

OpenStudy (anonymous):

welcome